### v1

@@ 3.1 Transferability and accuracy of the physical model
In order to test our model’s ability to generalize across different RNA systems, we evaluated our unified physics-based framework on a set of well-characterized RNA structures previously studied using DMS chemical probing [bohnNanoDMSMaPAllowsIsoformspecific2023]. This set spans a range of structural complexity, comprising five distinct systems: (i) bacterial RNase P (type A), (ii) hc16 ligase, (iii) Tetrahymena ribozyme, (iv) V. cholerae glycine riboswitch, and (v) HCV IRES. These sequences range from approximately 100 to 400 nucleotides in length. Reference structures for hc16 ligase, Tetrahymena ribozyme, and V. cholerae glycine riboswitch were obtained from cryo-EM-guided structure determination [kappelAcceleratedCryoEMguidedDetermination2020], and secondary structure was annotated from the 3D structure with Barnaba [bottaroBarnabaSoftwareAnalysis2019]. HCV IRES reference structure was manually transcribed from Extended Data Fig. 6 of Ref.  [bohnNanoDMSMaPAllowsIsoformspecific2023]. For bacterial RNase P (type A), although a crystal structure exists [torres-lariosCrystalStructureRNA2005], sequence differences between the crystallized construct and the chemically probed sequence preclude its use as a reference. These reference structures are expected to be more accurate with respect to the predictions done using a simple thermodynamic model. However, since they represent single structures, they do not report on conformational heterogeneity.
@@
Figure 2: (a) Cross-validation of physical parameters across five RNA systems. Each bar represents the per-datapoint loss obtained for the indicated system after training on three systems and testing on the remaining two. Yellow bars refer to results obtained when the corresponding system was excluded from the training set. (b) Normalized loss profiles for HCV IRES with soft constraints of varying magnitude applied. Red line: constraints from reference structure; light green line: constraints from minimum free energy (MFE) structure. Each curve represents an independent fit with physical parameters trained on different system triplets. The blue line corresponds to an improved fit achieved by optimizing the sequence-specific $\lambda_{i}$ parameters with MERGE-RNA. Stars mark are reused across panels to indicate the same ensemble. (c) Frobenius distance between base pairing probability (bpp) matrices, quantifying structural similarity between the HCV IRES reference structure and: MFE structure, bpp of the original thermodynamic ensemble, bpp obtained from the minimum of the red curve in panel (b), and bpp from the fully optimized model. (d) Arc plot visualization comparing the secondary structures of HCV IRES: reference experimental structure (red), ensemble obtained from the thermodynamics model (green), and ensemble obtained from MERGE-RNA (blue).
@@
Our first objective was to assess whether the physical parameters extracted from our model represent genuine physical properties of the experiment and are hence transferable to different systems, rather than being overfitted on a specific case. We performed comprehensive cross-validation fits, training the physical parameters of the model (all the parameters except the $\lambda_{i}$) on triplets of RNA systems and testing on the remaining pairs (Fig. 2a). Loss values (as defined in Equation 31) are then divided by the number of data points to enable direct comparison. The loss values obtained across all sets are consistent both in training and in test, demonstrating that our model generalises on multiple systems and successfully captures the underlying principles governing chemical probing experiments. Notably, physical parameters obtained through different minimizations might be different (see Table 1) but typically predict similar mutation profiles. One intuitive example of this is that by increasing the reference chemical potential ($\mu_{r}$), the predicted mutation rate will–on average–increase, but this can be compensated by a decrease of binding probability $p_{\text{bind}}$. This under-determination of parameters can arise when the experiments do not cover a wide enough range of probe concentrations, but it does not affect the model’s ability to predict reactivity profiles within the range of conditions upon which it was trained.
3.2 Importance of conformational heterogeneity in structured RNAs
@@
Having established the generalizability of the physical parameters, we next used our framework to explore the balance between structural accuracy and ensemble heterogeneity. Using HCV IRES as a representative example (analogous results for other systems in Supplementary Figs. 6,7,8), we varied a uniform soft constraint (Fig. 2b): for a given scalar magnitude (x-axis), every nucleotide annotated as paired in the chosen reference structure (experimental or minimum-free-energy) was assigned that value and all other positions were set to zero, with no position-specific tuning. Increasing this magnitude progressively biases the ensemble toward the targeted structure, smoothly interpolating between the original thermodynamic ensemble and a single constrained structure while keeping the physical parameters fixed.
@@
By interpolating between the original ensemble and its single most-likely structure (minimum-free-energy), we assess the role of conformational heterogeneity (light green line in Fig. 2b). In fact, favoring a single structure has the effect of progressively suppressing the population of alternative conformations that are likely present in solution and contribute to the observed mutation profiles. Interestingly, the loss function monotonically increases when the constraint is applied, degrading the agreement with experiment. On the other hand, incorporating knowledge from the reference experimental structure (red line in Fig. 2b) initially improves the model’s agreement with experimental data. However, even in this case, excessive constraint enforcement eventually degrades performance. This pattern indicates the existence of a “sweet spot” where the model partially incorporates reference structural information while preserving the conformational diversity necessary to explain the experimental measurements. Too strong a constraint toward any single structure—even the experimental reference—reduces the model’s ability to account for the heterogeneous structural ensemble that is captured in the experimental data. Overall, this analysis shows that a proper balance of information from the reference single experimental structure and ensemble diversity improves the capability of the model to reproduce chemical probing experiments.
3.3 Inference of optimal ensemble
@@
Building on these insights, instead of borrowing from the reference structure, we inferred the best sequence-specific soft constraints ($\lambda_{i}$ parameters) using our model. This resulted in substantially improved agreement with experimental data (blue line in Fig. 2b), as expected given that the experimental data is used to infer the soft constraints.
@@
To validate our model independently, we computed the Frobenius distance between the reference structure and base-pairing probability matrices obtained by the baseline ViennaRNA MFE, Vienna BPP, and those optimally guided by the reference structure (Fig. 2c). Our model achieves comparable accuracy for some systems and clear improvements for others (Figs. 6,7,8). Remarkably, our inference performs strongly even against ensembles obtained using uniform soft constraints that are reference-informed (red star on the red curve in Fig. 2b) yet our model does this without knowledge of the underlying experimental structure. This apparent paradox arises because our model fits an array of constraints individually—allowing for greater flexibility, whereas enforcing the reference structure modifies the magnitude of all constraints simultaneously. A visualization of these improvements is illustrated through arc plot representations in Fig. 2d, which demonstrate that our model-derived ensemble (blue) captures the conformational diversity while recapitulating the reference structure (red) more accurately than the standard thermodynamic ensemble (green). Similar improvements were observed across all tested RNA systems, confirming the broad applicability of our approach.
@@
In summary, with our model (1) the inferred physical parameters are generalisable across systems (Fig. 2a); (2) there is an optimal balance between structural accuracy and ensemble heterogeneity, and inference of sequence-specific corrections further improves agreement with experimental data (Fig. 2b); and (3) the resulting structural ensembles more accurately capture native conformations compared to standard thermodynamic approaches (Fig. 2c,d).
3.4 Application on a known case of functional structural rearrangement: cspA 5’ UTR
@@
We want to test our model’s ability to capture complex structural rearrangements in relevant biological systems, so we consider two DMS mutational profiles for cspA 5 UTR recorded at 10 and 37C [zhangStressResponseThat2018]. This sequence is known to act as a thermoswitch [giuliodoriCspAMRNAThermosensor2010], with different structures at the two reported temperatures. The two experiments have been recorded at different probing times (2 hours vs 10 minutes). Our approach has been designed to analyze homogeneous experiments and does not model the dependence of the observed mutation rate on the probing time. We hence normalize the dataset at 10C to the same average mutation rate observed at 37C. In addition, to avoid confusing the effect of the mutational profiles with the temperature-dependence of the thermodynamic model, we use the thermodynamic model set on the same intermediate temperature for every ensemble (23.5C). Finally, we note that these experimental datasets lack a control mutation profile obtained in absence of DMS, making the separation of the physical signal from the systematic bias more difficult for our model. Given the metioned limitations, we decided to perform a set of 6 independent fits, starting from different random initializations of the physical parameters. We report in Fig. 3 the ensemble predicted by the model that achieved the lowest loss, and the other fits are reported in Figs. 9-13 as Supplementary Information.
Figure 3(a,b) reports the ensembles predicted by MERGE-RNA, which are similar to the individual structures predicted in a previous work [zhangStressResponseThat2018]. Interestingly, the ensembles that we predict at 37C (10C) are very similar to those predicted in ref. [zhangStressResponseThat2018] in positions 70–170 (1–120), but differ in other regions. We note that our model reports a thermodynamic ensemble, where multiple competitive helices might appear with the proper population, where Ref. [zhangStressResponseThat2018] reports a single minimum-free-energy structure.
Overall, this analysis demonstrates that our model can capture complex structural rearrangements in biological systems.
3.5 Deconvolution of mixed structural states on synthetic data
@@
To test the model’s ability to predict the correct secondary-structure populations in heterogeneous ensembles, we generated synthetic datasets for an in-house designed bistable RNA carrying two mutually exclusive hairpins. We deliberately designed this construct so that we could also assay the very same sequence experimentally (discussed in the next section); accordingly, we appended primer binding sites at the 3 and 5 ends. Concretely, the final construct has the architecture
@@
where PBS1 and PBS2 are distinct primer-binding sites, A1 and A2 are identical copies of the same sequence, and B is complementary to A1/A2. In this design, A1 and A2 compete to pair with B, producing two mutually exclusive hairpins (helix 1: A1:B; helix 2: A2:B; as schematically visualised in Fig. 5c, see ”baseline prediction”). L1 and L2 are short linkers designed to act as tetraloops for the hairpins.
@@
We then trained our model on these synthetic datasets. Following the procedure established earlier, we first performed a joint fit of the physical parameters across all datasets, starting from random initializations. Subsequently, we optimized the soft constraints $\lambda_{i}$ for each dataset individually. The results, presented in Fig. 4, show that the model accurately recovers the expected base-pairing probabilities for both helices across all synthetic datasets. The method is robust to the presence of additional base pairs formed by the random primers. This serves as a proof-of-concept for the model’s ability to deconvolve mixed structural states from chemical probing data across a wide range of populations (Fig. 4c).
3.6 A putatively bistable sequence exhibits a heterogeneous ensemble with strand displacement
@@
To validate our model on experimental data, we collected DMS-MaP data on the putative bistable RNA construct described in the previous section. The data was collected at multiple DMS concentrations, 0 to 100 mM at steps of 20 mM, 2 replicas each. The obtained mutation profiles are visualised in Fig. 5a. We then trained our model on these experimental datasets, following the same procedure as in the previous section. Note that the sequence, as discussed earlier, was designed to adopt two mutually exclusive helices. We then applied our model to infer the structural ensembles from the data.
@@
The results of our model fitting are presented in the same panel as solid lines. Remarkably, the two prediction of MERGE-RNA are consistent with each other, regardless of the initial thermodynamic parameter used, signaling robustnees of the prediction with respect to the baseline parameters. In both cases, the model adjusts the ensemble of structure so that the resulting ensemble slightly favors helix 1, in accordance with the higher mutation rates observed in the experimental data (Fig. 5a).
@@
Crucially, the model does more than just correct the populations of the two helices. It also predicts a gradient in base-pairing probabilities, where, within regions A1 and A2, nucleotides closer to the central domain (B) are more likely to be paired. This is a direct consequence of a corresponding slope in the observed reactivities (Fig. 5a). In fact, the model captures not only the two fully-formed helices but also intermediate states where the central domain is partially paired with both the competing strands, as schematically illustrated in Fig. 5c. This mechanism, known as strand displacement (or strand invasion) [ratajczykControllingDNARNA2025], has been observed in biological systems and is consistent with the observed mutation profile. In fact, regions nearest the central domain are expected to pair first during displacement, and, accordingly, the mutation profile shows reduced reactivity closer to the loops (Fig. 5a). The presence of intermediate states where a fraction of each of the two helices is formed is expected if the two loops are sufficiently stable. The loop co-occupancy analysis in Table 2 shows that baseline ensembles rarely populate states with both loops formed simultaneously $(<0.2\%)$, whereas MERGE-RNA yields concurrent formation in 42–57% of structures.
This analysis underscores the importance of considering heterogeneous ensembles in RNA structure prediction, as they can capture dynamic processes such as strand displacement that are not accounted for in traditional Minimum Free Energy (MFE) approaches. Note that, since the potential strand displacement intermediates had a negligible population in the baseline thermodynamic model, this result would have very likely been missed by traditional approaches that enumerate and reweight only a limited set of suboptimal structures.
4 Discussion

### v2

@@ Results
We extracted physical parameters from our unified physics-based framework after training on a set of well-characterized RNA structures previously studied using DMS chemical probing [bohnNanoDMSMaPAllowsIsoformspecific2023]. This set spans a range of structural complexity, comprising five distinct systems: (i) bacterial RNase P (type A), (ii) hc16 ligase, (iii) Tetrahymena ribozyme, (iv) V. cholerae glycine riboswitch, and (v) HCV IRES. These sequences range from approximately 100 to 400 nucleotides in length. Reference structures for hc16 ligase, Tetrahymena ribozyme, and V. cholerae glycine riboswitch were obtained from cryo-EM-guided structure determination [kappelAcceleratedCryoEMguidedDetermination2020], and secondary structure was annotated from the 3D structure with Barnaba [bottaroBarnabaSoftwareAnalysis2019]. HCV IRES reference structure was manually transcribed from Extended Data Fig. 6 of Ref.  [bohnNanoDMSMaPAllowsIsoformspecific2023]. For bacterial RNase P (type A), although a crystal structure exists [torres-lariosCrystalStructureRNA2005], sequence differences between the crystallized construct and the chemically probed sequence preclude its use as a reference. These reference structures are expected to be more accurate with respect to the predictions done using a simple thermodynamic model. However, since they represent single structures, they do not report on conformational heterogeneity.
@@
Figure 2: (a) Cross-validation of physical parameters across five RNA systems. Each bar represents the per-datapoint loss obtained for the indicated system after training on three systems and testing on the remaining two. Yellow bars refer to results obtained when the corresponding system was excluded from the training set. (b) Normalized loss profiles for HCV IRES with soft constraints of varying magnitude applied. Red line: constraints from reference structure; light green line: constraints from minimum free energy (MFE) structure. Each curve represents an independent fit with physical parameters trained on different system triplets. The blue line corresponds to an improved fit achieved by optimizing the sequence-specific $\lambda_{i}$ parameters with MERGE-RNA. Stars are reused across panels to indicate the same ensemble. (c) Frobenius distance divided by the squared number of nucleotides between base pairing probability (bpp) matrices, quantifying structural similarity between the HCV IRES reference structure and: MFE structure, bpp of the original thermodynamic ensemble, bpp obtained from the minimum of the red curve in panel (b), and bpp from the fully optimized model. (d) Mutation rates at 57mM after background subtraction for adenine and cytosine in HCV IRES, grouped by pairing status in the reference structure. Each point is colored according to the pairing probability predicted by MERGE-RNA. Strong disagreements (positions where MERGE-RNA predicts pairing probability below 25% or above 75% while the reference indicates the opposite) are highlighted.
Figure 3: Arc plot visualization comparing the secondary structures of HCV IRES: reference experimental structure (red), ensemble obtained from the thermodynamics model (green), and ensemble obtained from MERGE-RNA (blue).
@@
Our first objective was to assess whether the physical parameters extracted from our model represent genuine physical properties of the experiment and are hence transferable to different systems, rather than being overfitted on a specific case. We performed comprehensive cross-validation fits, training the physical parameters of the model (all the parameters except the $\lambda_{i}$) on triplets of RNA systems and testing on the remaining pairs (Fig. 2a). Loss values (as defined in Equation 31) are then divided by the number of data points to enable direct comparison. The loss values obtained across all sets are consistent both in training and in test, demonstrating that our model generalizes on multiple systems and successfully captures the underlying principles governing chemical probing experiments. Notably, physical parameters obtained through different minimizations might be different (see Table S1) but typically predict similar mutation profiles. One intuitive example of this is that by increasing the reference chemical potential ($\mu_{r}$), the predicted mutation rate will–on average–increase, but this can be compensated by a decrease of binding probability $p_{\text{bind}}$. This under-determination of parameters can arise when the experiments do not cover a wide enough range of probe concentrations, but it does not affect the model’s ability to predict reactivity profiles within the range of conditions upon which it was trained.
Importance of conformational heterogeneity in structured RNAs
@@
Having established the generalizability of the physical parameters, we next used our framework to explore the balance between structural accuracy and ensemble heterogeneity. Using HCV IRES as a representative example (analogous results for other systems in Supplementary Figs. S1,S2,S3), we varied a uniform soft constraint (Fig. 2b): for a given scalar magnitude (x-axis), every nucleotide annotated as paired in the chosen reference structure (experimental or minimum-free-energy) was assigned that value and all other positions were set to zero, with no position-specific tuning. Increasing this magnitude progressively biases the ensemble toward the targeted structure, smoothly interpolating between the original thermodynamic ensemble and a single constrained structure while keeping the physical parameters fixed.
@@
By interpolating between the original ensemble and its single most-likely structure (minimum-free-energy), we assess the impact of reducing ensemble heterogeneity on the agreement with experiment (light green line in Fig. 2b). In fact, favoring a single structure has the effect of progressively suppressing the population of alternative conformations that are likely present in solution and contribute to the observed mutation profiles. Interestingly, the loss function monotonically increases when the constraint is applied, degrading the agreement with experiment. On the other hand, incorporating knowledge from the reference experimental structure (red line in Fig. 2b) initially improves the model’s agreement with experimental data. However, even in this case, excessive constraint enforcement eventually degrades performance. This pattern indicates the existence of a “sweet spot” where the model partially incorporates reference structural information while preserving the conformational diversity necessary to explain the experimental measurements. Too strong a constraint toward any single structure—even the experimental reference—reduces the model’s ability to account for the heterogeneous structural ensemble that is captured in the experimental data. Overall, this analysis shows that a proper balance of information from the reference single experimental structure and ensemble diversity improves the capability of the model to reproduce chemical probing experiments.
Inference of optimal ensemble
@@
Building on these insights, instead of borrowing from the reference structure, we inferred the best sequence-specific soft constraints ($\lambda_{i}$ parameters) using our model. This resulted in substantially improved agreement with experimental data (blue line in Fig. 2b), as expected given that the experimental data are used to infer the soft constraints.
@@
To validate our model independently, we computed the Frobenius distance between the reference structure and base-pairing probability matrices obtained by the baseline ViennaRNA MFE, Vienna BPP, and those optimally guided by the reference structure (Fig. 2c). Our model achieves comparable accuracy for some systems and clear improvements for others (Figs. S1–S3 in SI Appendix). Remarkably, our inference performs strongly even against ensembles obtained using uniform soft constraints that are reference-informed (red star on the red curve in Fig. 2b) yet our model does this without knowledge of the underlying experimental structure. This performance is likely due to our model fitting each nucleotide’s soft constraint ($\lambda_{i}$) individually, whereas the reference-informed approach applies a uniform constraint to all paired positions. Arc plots (Fig. 3) illustrate how this MERGE-RNA’s individualized fitting preserves ensemble diversity (blue) while recapitulating the reference structure (red) more accurately than the standard thermodynamic ensemble (green). Similar improvements were observed across all tested RNA systems, confirming the broad applicability of our approach.
MERGE-RNA reconciles dynamic probing data with static reference structures
Having established structural accuracy by comparison with experimentally resolved reference structures, we now examine where our predictions diverge from the static reference. Since our model predicts a dynamic ensemble rather than a single conformation, divergence is expected and informative: chemical probing experiments inherently report on the dynamic ensemble of the RNA molecule, which a static reference structure alone cannot fully capture. In Figure 2d we report mutation rates after control subtraction for adenine and cytosine, grouped by pairing status in the reference structure. As expected, paired positions show, on average, lower mutation rates than unpaired ones. Each point is colored according to the pairing status predicted by MERGE-RNA. The model’s predictions correlate well with mutation rates, demonstrating its capability to infer and model pairing dynamics from probing data. It is instructive to examine positions where our predictions disagree with the reference structure.
When mutation rates are consistent with the reference structure, MERGE-RNA’s predictions align with it. For positions with intermediate mutation rates, MERGE-RNA predicts intermediate pairing probabilities, suggesting conformational heterogeneity in solution. This is particularly informative: a static reference structure classifies each position as either paired or unpaired, and thus cannot account for intermediate reactivity. MERGE-RNA bridges this knowledge gap by proposing alternative conformations that together explain the observed mutation rates. Where probing data and reference structure are in contrast, MERGE-RNA captures new pairings absent from the reference but supported by probing, or identifies reference-annotated pairs that show elevated mutation rates, suggesting they are frequently unpaired in the dynamic ensemble.
The number of “strong disagreements sites” (positions where MERGE-RNA predicts pairing probability below 25% or above 75% while the reference indicates the opposite for A and C nucleotides, highlighted in Fig. 2d) is modest, ranging from 0 to 13 across the systems studied here. Detailed analysis (SI Appendix Figs. S4S7) reveals that most correspond to small corrections of existing structural elements or formation/dismantling of short motifs (e.g., 3–4 bp stems), indicating localized inconsistencies rather than wide structural disagreements. We note that some discrepancies may also arise from in vitro versus in-cell conditions, protein binding, errors in reference structure annotation, or other artifacts.
@@
In summary, with our model (1) the inferred physical parameters are generalisable across systems (Fig. 2a); (2) we show that there is an optimal balance between structural accuracy and ensemble heterogeneity, and inference of sequence-specific corrections further improves agreement with experimental data (Fig. 2b); (3) the resulting structural ensembles more accurately capture native conformations compared to standard thermodynamic approaches (Fig. 2c); (4) comparing MERGE-RNA’s results to the reference structures reveals that agreement reflects probing data consistency (Fig. 2d). In cases where reference models are inconsistent with experiment, MERGE-RNA goes beyond the static reference by identifying localized discrepancies and proposing alternative conformations that better explain the data. (5) these results are consistent across all tested RNA systems (SI Appendix Figs. S1S3).
Application on a known case of functional structural rearrangement: cspA 5’ UTR
@@
We next tested our model’s ability to capture complex structural rearrangements in relevant biological systems by considering two DMS mutational profiles for cspA 5 UTR recorded at 10 and 37C [zhangStressResponseThat2018]. This sequence acts as a thermoswitch [giuliodoriCspAMRNAThermosensor2010], with different structures at the two reported temperatures. The two experiments have been recorded at different probing times (2 hours vs 10 minutes). Our approach has been designed to analyze homogeneous experiments and does not model the dependence of the observed mutation rate on the probing time. We hence normalize the dataset at 10C to the same average mutation rate observed at 37C. In addition, to avoid confusing the effect of the mutational profiles with the temperature-dependence of the thermodynamic model, we use the thermodynamic model set on the same intermediate temperature for every ensemble (23.5C). Finally, we note that these experimental datasets lack a control mutation profile obtained in absence of DMS, making the separation of the physical signal from the systematic bias more difficult for our model. Given the mentioned limitations, we decided to perform a set of 6 independent fits, starting from different random initializations of the physical parameters. We report in Fig. 4 the ensemble predicted by the model that achieved the lowest loss, and the other fits are reported in the SI Appendix (Figs S8S12).
Figure 4(a,b) shows the structural ensembles predicted by MERGE-RNA, which largely recapitulate the individual structures reported by Zhang et al. [zhangStressResponseThat2018]. At 10C and 37C, our ensembles closely match their predictions in positions 1–120 and 70–170, respectively, but exhibit notable differences in other regions, particularly in segments where alternative helices compete. Unlike Ref. [zhangStressResponseThat2018], which reports a single minimum-free-energy structure, MERGE-RNA predicts a thermodynamic ensemble that accounts for multiple competitive helices and their relative populations. Overall, this analysis demonstrates that our model can capture complex structural rearrangements in biological systems.
Deconvolution of mixed structural states on synthetic data
@@
To rigorously test the model’s ability to correctly predict secondary-structure populations in heterogeneous ensembles, we generated synthetic datasets for an in-house designed bistable RNA carrying two mutually exclusive hairpins. We deliberately designed this construct so that we could also assay the very same sequence experimentally (discussed in the next section); accordingly, we appended primer binding sites at the 3 and 5 ends. Concretely, the final construct has the architecture
@@
where PBS1 and PBS2 are distinct primer-binding sites, A1 and A2 are identical copies of the same sequence, and B is complementary to A1/A2. In this design, A1 and A2 compete to pair with B, producing two mutually exclusive hairpins (helix 1: A1:B; helix 2: A2:B; as schematically visualized in Fig. 6c, see ”baseline prediction”). L1 and L2 are short linkers designed to act as tetraloops for the hairpins.
@@
We then trained our model on these synthetic datasets. Following the procedure established earlier, we first performed a joint fit of the physical parameters across all datasets, starting from random initializations. Subsequently, we optimized the soft constraints $\lambda_{i}$ for each dataset individually. The results, presented in Fig. 5, show that the model accurately recovers the expected base-pairing probabilities for both helices across all synthetic datasets. The method is robust to the presence of additional base pairs formed by the random primer binding sites. Overall, these synthetic data tests validate the model’s ability to resolve heterogeneous RNA ensembles and recover helix populations consistent with the known ground truth (Fig. 5c).
A putatively bistable sequence exhibits a heterogeneous ensemble with strand displacement
@@
To validate our model on experimental data, we collected DMS-MaP data on the putative bistable RNA construct described in the previous section. The data were collected at multiple DMS concentrations, 0 to 100 mM at steps of 20 mM, 2 replicas each. The obtained mutation profiles are visualized in Fig. 6a. We then trained our model on these experimental datasets, following the same procedure as in the previous section. Note that the sequence, as discussed earlier, was designed to adopt two mutually exclusive helices. We then applied our model to infer the structural ensembles from the data.
@@
The results of our model fitting are presented in the same panel as solid lines. Remarkably, the two predictions of MERGE-RNA are consistent with each other, regardless of the initial thermodynamic parameter used, signaling robustness of the prediction with respect to the baseline parameters. In both cases, the model adjusts the ensemble of structures so that the resulting ensemble slightly favors helix 1, in accordance with the lower mutation rates observed in the experimental data (Fig. 6a).
@@
Crucially, MERGE-RNA does more than adjust the populations of the two competing helices. It reveals a gradient in base-pairing probabilities within regions A1 and A2, where nucleotides closer to the central domain (B) are more likely to be paired, reflecting the slope observed in the experimental mutation profiles (Fig. 6a). Beyond the two fully-formed helices, the model resolves intermediate states in which B is partially paired with both A1 and A2, as schematically illustrated in Fig. 6c. This mechanism, known as strand displacement (or strand invasion) [ratajczykControllingDNARNA2025], has been observed in biological systems and is consistent with the observed mutation profile: regions nearest the central domain are expected to pair first during displacement, and accordingly show reduced reactivity closer to the loops (Fig. 6a). The presence of such intermediate states is expected if the two loops are sufficiently stable. Analysis of loop co-occupancy (Table S2) demonstrates capture of intermediate states missed by standard models. While baseline ensembles rarely allow both loops to form (<0.2<0.2%), the model predicts simultaneous pairing in 42–57% of structures, consistent with the observed mutation profiles and reflecting the strand-displacement mechanism in which the central domain transiently pairs with both competing helices. Overall, these results demonstrate MERGE-RNA’s ability to capture a continuous spectrum of low-population intermediates that standard thermodynamic models fail to predict, highlighting its ability to reveal dynamic structural ensembles inaccessible to conventional approaches, independently of the baseline thermodynamic model.
Discussion